OpenStudy (anonymous):

the mean of 100 numerical observations is 35.73. What is the value of the sum of all 100 numbers?

OpenStudy (lalaly):

\[mean=\frac{\sum x}{n}\] \[35.73=\frac{ \sum x}{100}\]\[\sum x =3573\]

OpenStudy (anonymous):

so the answer is 3573?

OpenStudy (lalaly):

yeah because the mean is sum of the numbers divided by the size of the collection.

OpenStudy (anonymous):

thank-you lana!!

OpenStudy (lalaly):

I hope it helped:)

OpenStudy (anonymous):

it did ^_^

OpenStudy (lalaly):

:D:D